OpenStudy (anonymous):

The volume of a sphere is 3,000pi m^3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (anonymous):

volume = 4/3 pi r^3 find r and use that to calculate surface area

OpenStudy (anonymous):

V = (4/3) * pi * r^3 3000pi = (4/3) * pi r^3 r^3 = 2250 r = 2250^(1/3) SA = 4*pi*r^2 SA = 4*pi*(2250^(1/3))^2 SA = 4*pi*2250^(2/3) Simplify as needed.